40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Aug 22 02:40:12 2013, Jenne, Update, LSC, DRMI Locked for 1+ minute!!!!!! DRMI_First_Lock_1plusMinute_21Aug2013.png
    Reply  Thu Aug 22 07:57:57 2013, Lisa, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
       Reply  Thu Aug 22 13:03:40 2013, Jenne, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Thu Aug 22 10:20:32 2013, kiwamu, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Thu Aug 22 13:20:54 2013, Koji, Update, LSC, DRMI Locked for 1+ minute!!!!!! 
    Reply  Sat Aug 24 00:11:07 2013, Koji, Update, LSC, DRMI Locked with improved lock streatches Screenshot-Untitled_Window.png
       Reply  Sun Aug 25 06:07:11 2013, rana, Update, LSC, DRMI Locked with improved lock streatches 
Message ID: 9060     Entry time: Sat Aug 24 00:11:07 2013     In reply to: 9049     Reply to this: 9061
Author: Koji 
Type: Update 
Category: LSC 
Subject: DRMI Locked with improved lock streatches 

Friday night locking

Much more stable DRMI lock was achieved, partly thanks to the Friday-night quiet seismic,
and partly because of the improved servo gain and LF boosts


55MHz thru-put

I wanted to confirm the enhancement of the 110MHz signal at the AS port.

As the AS110 PD is placed in the CCD path, there is nothing visible with PRMI.
The Thorlabs PD was moved to the main AS path. Now the AS110 PD is receiving 50% of the power.
 

With PRMI 110MHz peak was -30dBm (As it was fluctuating, anything more precise number did not make sense)
When the DRMI was locked, the peak was enhanced to 0dBm.

The 2f signal comes from the beat between the sidebands.
Thus the amplitude of the intensity is proportional to the power of the sidebands (assuming the +1 and -1 order sidebands have the same amplitude)
-30dBm -> 0dBm means 31.6 times amplitude of the intensity. Therefore the amplitude transmission of the sidebands is 5.6 times more. (Is this true?)

According to the wiki, the AS port thru-put (i.e. power transmission) for the 55MHz sideband is 0.0026 and 0.43 for PRMI and DRMI respectively.
This corresponds to the amplitude difference of ~13. So we still have only half of the sidebands leaking out from the IFO. This could be attributed
to both the smaller PR gain and SR gain.


Locking setup

Same as the one Jenne used the other day. Later I engaged several additional triggers.
The following is the trigger setting I used

MICH: Delay 2 sec, FM1/FM2/FM3/FM6/FM7

PRCL: Delay 0.5 sec, FM2/FM3/FM6

SRCL: Delay 5 sec, FM1/FM2/FM3/FM6

SRCL FM1 was modified from +3dB to +6dB


Lock stability

Once lock is acquired, it lasts tens of minutes. (see the attached striptool chart.)
Even the lock is lost, it reacquires quickly.

The videos to show the lock acquisition and the in-lock stability are attached below.
The AS port beam is very round. It is not so shaky, but some yaw motion is visible.
The mode at the AS port is defined by the SRM, putting a QPD at the AS port would help to
stabilize the spot.


IFO state upon leaving

I left the 40m with the arms aligned, PRM and SRM slightly misaligned, and LSC setting is for the DRMI locking.


TO DO

- AS110I/Q for triggering

- PRCL/MICH/SRCL normalization

- We should resurrect the IFO config scripts.

- Remove BS->SRCL actuation coupling

- Handing off to 3f signals (preparation for the full lock)

- Improve ALS stability

- SRM ASC: AS QPD for SRM control


Lock Acquisition Video
UL (REFL) / UR (POP)
LL (AS) / LR (PRM Face)

 
In-lock video
UL (REFL) / UR (POP)
LL (AS) / LR (PRM Face)

 

Attachment 1: Screenshot-Untitled_Window.png  26 kB  Uploaded Sat Aug 24 01:17:18 2013  | Hide | Hide all
Screenshot-Untitled_Window.png
ELOG V3.1.3-